The rms voltage across a 2.50 uF capacitance is 25.0 V if the rms currents for the frequencies 100 Hz, 1000 Hz and 5000 Hz are I1, I2 and I3 respectively, then one can say

I1 > I2 >I3

I1 > I3 >I2

I3 > I2 >I1

I3 > I1 >I2
Answer #1

The capacitive reactance of the capacitor is

XC =1 / 2\pifC

Here, frequency is f and capacitance is C.

Now, the rms current flows through the capacitor is

Irms = Vrms/ XC

         = Vrms / (1 / 2\pifC)

      = Vrms 2 \pifC

For I1,

I1 = (25.0 V) 2 \pi (100 Hz) (2.50 uF)(10-6 F / 1 uF)

     = 0.03925 A

For I2,

I2 = (25.0 V) 2 \pi (1000 Hz) (2.50 uF)(10-6 F / 1 uF)

     = 0.3925 A

For I3,

I3 = (25.0 V) 2 \pi (1000 Hz) (2.50 uF)(10-6 F / 1 uF)

     = 1.9625 A

Therefore, the rms current follows I3 > I2 > I1.